A chemist has a beaker of a 3% acid solution and a beaker of a 7% solution. He needs to make 75ml of a 4% acid solution.
Amount of solution (ml): x 3% solution + y 7% solution = ---------
Amount of Acid (ml) : _______x + ______y = 0.04(75)
Use the information in the table to write a system of linear equations.
Solve the system of equations to find how much he will use from each beaker.

Answer by mananth(16949)

A chemist has a beaker of a 3% acid solution and a beaker of a 7% solution. He needs to make 75ml of a 4% acid solution.
Amount of solution (ml): x 3% solution + y 7% solution = 75---------
Amount of Acid (ml) : __0.03_____x + __0.07____y = 0.04(75)
Use the information in the table to write a system of linear equations.
Solve the system of equations to find how much he will use from each beaker.
x+y=75
0.03x+0.07y=0.04*75
Multiply the equation by 100
3x+7y= 300
3x+3y-3x-7y= 225-300
-4y=-75
y= 18.75 ml
x=75-18.75
x=56.25 ml